OpenStudy (anonymous):

Give the following three vectors u=2i-j+k, v=i+j+2k, w=2i-j+4k. Determine the scalars such that w-gu-hv is perpendicular to both u and v.

OpenStudy (amistre64):

determine scalars g and h such that .... = uxv

OpenStudy (turingtest):

^better than my explanation :D

OpenStudy (amistre64):

i was hoping your explanation would fill in some of the details :)

OpenStudy (anonymous):

could you eloborate :)

OpenStudy (anonymous):

i just need something to get me started

OpenStudy (amistre64):

keep g and h generac as you combine the vectors; and then equate it to uxv to solve for the scalars

OpenStudy (turingtest):

\[\vec w-g\vec u-h\vec v=...\]write out the expression for the components and call the resulting vector r we know that two vectors are perpendicular if their dot product is zero\[\vec r\cdot\vec u=\vec r\cdot\vec v=0\]solve for the scalars

OpenStudy (anonymous):

hmm, okay thanks i get what you mean

OpenStudy (phi):

are you learning about Gram-Schmidt?

OpenStudy (phi):

find the projection, and subtract it off is the basic idea.

OpenStudy (anonymous):

oh no, just calc 3 chap on cross product.

OpenStudy (anonymous):

i will good Gram-Shmidt, sure seems easier

OpenStudy (anonymous):

google*

OpenStudy (amistre64):

-gu=<-2g,1g,-1g>, -hv=<-1h,-1h,-2h>, w=<2,-1,4> w=< 2 ,-1 , 4 > -gu=<-2g,1g,-1g> -hv=<-1h,-1h,-2h> ------------------- <2-2g-h, g-h-1, 4-g-2h> = uxv

OpenStudy (amistre64):

uxv: x 2 1 x=-2-1 y -1 1 -y=(4-1) z 1 2 z = 2+1 uxv = <-3,-3,3>

OpenStudy (phi):

yes, but they expect you to use the cross product, see amistre

OpenStudy (amistre64):

compare parts :) -3 = 2-2g-h -5 = -2g-h h = 5-2g ; see below, since g=1, h=3 -3 = g-h-1 -3 = g-(5-2g)-1 -2 = g-5+2g 3 = 3g; g=1 lets see if this makes the final result work out, crosses fingers 3 = 4-g-2h 3 = 4 -1-2(3) 3 = 4-1-6 3=-3 hmmm its close so im sure its a typo in there someplace

OpenStudy (anonymous):

very cool, thanks now let me go ahead and try it myself,so i can be ready for this exam i got next week

OpenStudy (phi):

you'll get it.

OpenStudy (amistre64):

the concepts good i believe so just keep an eye on the mathing lol